# From Restless to Contextual: A Thresholding Bandit Reformulation for Finite-horizon Improvement

### v4

@@ 3 Optimality of BT-CMAB Reduction
Assuming all agents are homogeneous, i.e., $P^{m}=P^{m^{\prime}},R^{m}=R^{m^{\prime}},\forall m,m^{\prime}\in[M]$, $\bm{\pi}^{\text{greedy}}=\bm{\pi}^{*}$ for any $0\leq B\leq M<\infty$.
Theorem 3.1 establishes that our reference policy $\bm{\pi}^{b}$—which corresponds to the greedy policy $\bm{\pi}^{\text{greedy}}$ with appropriate threshold selection—achieves non-asymptotic optimality for homogeneous, 2-state finite-horizon restless bandits. This provides stronger guarantees than prior reference policies like WIP or LP-based methods (Hong et al., 2024; Whittle, 1988; Zhang & Frazier, 2021), which only offer asymptotic optimality for infinite-horizon settings under homogeneous agents.
We emphasize that while our analysis builds on the practical two-state assumption commonly used in recent work (Liang et al., 2024; Killian et al., 2023; Wang et al., 2023a), proving Theorem 3.1 required developing novel analytical techniques. Existing tools were designed for value function-based policies (Eq. (E.11)) in infinite-horizon settings (Whittle, 1988; Hong et al., 2024), making them unsuitable for analyzing our immediate-benefit greedy policy in finite-horizon contexts.
@@
Theorem 3.1 does not extend to heterogeneous agents (Appendix C) or multi-state systems (Appendix D). Our proof technique—which reduces Bellman optimality to conditions on value functions—relies critically on the two-state property that transition dynamics can be fully characterized by the probability of transitioning to state 1. This characterization breaks down in multi-state systems, where transitions require tracking probabilities to multiple states. While we conjecture that appropriately designed greedy policies may achieve constant-factor approximations even in these more general settings, such extensions are beyond the scope of this paper. Our primary contribution lies in identifying and formalizing the methodology of reformulating restless bandits to enable fast-converging algorithms (Sections 2.2-2.4), for which the two-state homogeneous setting provides a sufficiently rich foundation.
Before we end this section, we outline a proof sketch for Theorem 3.1. The proof consists of two steps: 1) We reformulate the problem as an MDP where the state space is the Cartesian product of the individual state space, i.e., $\mathcal{S}^{M}$; the action space is the subset of $\{0,1\}^{M}$ that satisfies the budget constraint, and the transition matrix is the product of $P^{m}$ as all agents are independent. 2) We show that the proposed greedy policy $\pi_{t}^{\text{greedy}}$ satisfies the Bellman equation (Eq. (E.21)) for all $t\leq T$ by using backward induction. We prove the induction hypothesis by carefully controlling the difference of the value function (Eq. (E.11)).
4 Budgeted Thresholding Contextual Multi-Armed Bandit
@@
13:   $\forall a\in\mathcal{A}:\hat{\phi}^{m}(s_{t}^{m},a)=\hat{\phi}^{m}(s_{t}^{m},a)+r^{m}_{t+1}\mathbbm{1}\left\{a_{t+1}^{m}=a\right\}$.
14:   $\forall a\in\mathcal{A}:N_{s,m,a}^{(t+1)}=N_{s,m,a}^{(t)}+\mathbbm{1}\left\{a_{t+1}^{m}=a\right\}\mathbbm{1}\left\{s=s_{t}^{m}\right\}$
15:   $N_{s,m}^{(t+1)}=N_{s,m}^{(t)}+\mathbbm{1}\{s=s_{t}^{m}\}$
@@
Before concluding this section, we emphasize that Alg. 1 differs from Auer et al. (2002). As we address a thresholding contextual bandit with MDP-governed contexts instead of MAB, we tailor $\epsilon_{t}$ to decay with the minimum state occupancy time rather than the current timestep, ensuring sufficient exploration across states.
@@ 4.1 Theoretical Guarantee
In this subsection, we establish that Algorithm 1 achieves a $\mathcal{O}(\sqrt{T})$ regret upper bound of under Assumps. 4.1 and 4.2. Assump. 4.1 ensures strictly positive transition probabilities (even if arbitrarily small), while Assump 4.2 guarantees the existence of a policy with sufficient good agent-state pairs at each timestep.
@@
Assump. 4.2 serves primarily to simplify our theoretical analysis by ensuring problem instances are well-behaved. This condition guarantees that an oracle policy exists which can spend its full budget of $B$ actions per round without being forced to assign actions with negative expected returns. In practice, this assumption is not restrictive. It can be satisfied for any problem instance by carefully choosing the threshold parameter $\gamma$; for example, setting $\gamma=$-$\infty$ guarantees the condition is met trivially. Thus, $\gamma$ serves as a tunable parameter to be optimized for overall performance, a point we discuss further in Remark 4.5.
When combined with Assump. 4.1, we can derive a stronger guarantee (Lemma 4.3) that is inherently assumed by any algorithm required to spend the full budget $B$ at each round. This is common in prior RB algorithms (Wang et al., 2020, 2023b). Let $\mathcal{G}:=\{(m,s):m\in[M],s\in\mathcal{S},I^{m}(s)\geq\gamma\}$ denote the set of good agent-state pairs. Lemma 4.3 ensures that for any state realization $\bm{s}\in\mathcal{S}^{M}$, there are always at least $B$ good agent-state pairs.
@@
Lemma 4.3 (proof in Appendix F.1), despite its strong guarantee, is derived from two relatively weak assumptions. This result is critical for two reasons. First, it ensures that any algorithm required to spend a full budget of $B$ actions is never forced to assign actions with negative expected returns, thereby avoiding linear regret. Second, and more profoundly, it guarantees that although selecting a suboptimal agent may affect the future state distribution, this effect does not accumulate over time because sufficient good agents are always available. This stability ensures that a policy achieving zero per-step regret is attainable, forming the foundation for our algorithm’s performance.
Let $\Delta_{s}^{m}:=|\gamma-I^{m}(s)|$. Let $\left(m_{\min},s_{\min}\right)$ be $\arg\min_{(m,s)\in\mathcal{G}}\left\{I^{m}(s)-\gamma\right\}$. And let $\Delta:=\Delta^{m_{\min}}_{s_{\min}}$. Next, we present our main theorem on the regret upper bound, which measures the speed at which Algorithm 1 is converging in a finite horizon:
@@
Assume $r_{t}^{m}$ is 1-subgaussian random variable for all $t\in[T]$. Let $D:=\max\left\{\frac{8M}{B},\frac{8M}{M-B},2\right\},p=1,\eta=\Delta$. The regret, $\mathcal{R}(\pi)$ , satisfies:
@@
Before concluding this section, we provide a proof sketch for Theorem 4.4. The argument proceeds in three steps. 1) We first show that analyzing the regret is equivalent to analyzing the expected number of times interventions are given to bad agents. 2) We then prove that this expected number can be bounded in terms of either the overestimation of the incremental reward of bad agents or the underestimation of the incremental reward of good agents. Moreover, we establish that at each time step, the number of interventions and non-interventions assigned to agents are both lower bounded. 3) Finally, we apply concentration inequality to obtain the desired result. Importantly, stationarity of the reward is not required here, since the concentration inequality we employ (Lemma F.2) holds without such an assumption.
@@
In our problem, there is an inherent trade-off between satisfying Assump. 4.2 and maximizing cumulative reward. While our reformulation focuses on the threshold identification problem, maximizing cumulative reward remains our ultimate goal. Specifically, a smaller threshold $\gamma$ makes it easier to satisfy Assump. 4.2, as more state-agent pairs qualify as good. However, a low $\gamma$ also makes any arm above the threshold optimal, potentially leading to lower rewards.
@@ 5 Experiments
Figure 1: Average (over time) cumulative reward for budgets $B=5$ (left), $10$ (middle), 2020 (right), averaged over 50 instances with 13 repetitions each. Top: noiseless reward; bottom: noisy reward.
@@
Setup   We set $M=50,\mathcal{S}=\{0,1\}$, and compare the results under three different budget settings: $B=5,10,20$. The expected reward is set to be $R^{m}(s)=s$ and randomly generate 50 instances with different transition matrices. We also consider the setting where $\mathcal{S}=[10]$ and the results are deferred to Appendix G.2.
In Figure 1, each line is averaged over $50$ instances, each with $13$ repetitions. We consider two reward settings: 1) a noiseless case, and 2) a noisy setting where the rewards at state $s\in\mathcal{S}$ follow a normal distribution with mean $R^{m}(s)$ and variance 1. We note that we did not finetune our hyperparameters. Specifically, we test two $\gamma$ and choose the best one and our algorithm is robust across different $\eta$ (Appendices G.2,G.3).
@@
Discussions   We establish greedy policy optimality in finite-horizon RBs and reformulate online RB as a BT-CMAB with a learning algorithm. Our approach leverages the insight that fast convergence to a suboptimal policy can outperform slow convergence to a better one in finite horizons. Future work includes extending to multi-state via careful reward design in BT-CMAB.
@@ Appendix B Whittle’s Index
$Q^{m}_{\lambda}(s,a)+\beta^{*}_{m}$ $=R^{m}(s)-\lambda\mathbbm{1}\{a=1\}$
@@
$\beta^{*}_{m}:=\max~\liminf_{T\to\infty}\frac{1}{T}\mathbb{E}\left[\sum_{t=0}^{T}R^{m}\left(s_{t}^{m}\right)-\lambda\mathbbm{1}\{a=1\}\mid s_{0}^{m}\right].$
@@
Theorem B.1 (Theorem 2, Weber & Weiss 1990).
@@
$Q^{m}_{\lambda}(s,a)$ $=R^{m}(s)-\lambda\mathbbm{1}\{a=1\}$
@@ Appendix E Technical Details In Section 3
$=\frac{I(0)-I(1)}{R(1)-R(0)}\left(F_{\bm{a}_{h+1},h+1}\left((1,0)\right)-F_{\bm{a}_{h+1},h+1}\left((0,0)\right)\right)-\frac{\left(q^{1}_{11}I(0)-q^{1}_{01}I(1)\right)}{R(1)-R(0)}\left(F_{\bm{a}_{h+1},h+1}\left((1,0)\right)-F_{\bm{a}_{h+1},h+1}\left((0,0)\right)\right)$
@@
$=\frac{I(0)-I(1)}{R(1)-R(0)}\left(F_{\bm{a}_{h+1},h+1}\left((1,0)\right)-F_{\bm{a}_{h+1},h+1}\left((0,0)\right)\right)-\frac{\left(q^{1}_{11}I(0)-q^{1}_{01}I(1)\right)}{R(1)-R(0)}\left(F_{\bm{a}_{h+1},h+1}\left((1,0)\right)-F_{\bm{a}_{h+1},h+1}\left((0,0)\right)\right)$

### v5

@@ 3 Optimality of BT-CMAB Reduction
Assuming all agents are homogeneous, i.e., $P^{m}=P^{m^{\prime}},R^{m}=R^{m^{\prime}},\forall m,m^{\prime}\in[M]$ and $\mathcal{S}=\{0,1\}$, $\bm{\pi}^{\text{greedy}}=\bm{\pi}^{*}$ for any $0\leq B\leq M<\infty$.
Theorem 3.1 establishes that our reference policy $\bm{\pi}^{b}$—which corresponds to the greedy policy $\bm{\pi}^{\text{greedy}}$ with appropriate threshold selection—achieves non-asymptotic optimality for homogeneous, 2-state finite-horizon restless bandits. This provides stronger guarantees than prior reference policies like WIP or LP-based methods (Hong et al., 2024; Whittle, 1988; Zhang and Frazier, 2021), which only offer asymptotic optimality for infinite-horizon settings under homogeneous agents.
While our analysis builds on the practical two-state assumption commonly used in recent work (Liang et al., 2024; Killian et al., 2023; Wang et al., 2023a), proving Theorem 3.1 required developing novel analytical techniques. Existing tools are designed for analyzing value function-based policies (Eq. (E.11)) in infinite-horizon settings (Whittle, 1988; Hong et al., 2024), making them unsuitable for our problem setup.
@@
Theorem 3.1 does not extend to heterogeneous agents (Appendix C) or multi-state systems (Appendix D). Our proof technique—which reduces Bellman optimality to conditions on value functions—relies critically on the two-state property that transition dynamics can be fully characterized by the probability of transitioning to state 1. This characterization breaks down in multi-state systems, where transitions require tracking probabilities to multiple states. While we conjecture that appropriately designed greedy policies may achieve constant-factor approximations even in these more general settings, such extensions are beyond the scope of this paper.
Although $\bm{\pi}^{b}$ is only guaranteed to be optimal for homogeneous agents and a high threshold, our BT-CMAB reduction is valid for heterogeneous agents and an arbitrary threshold $\gamma$. As mentioned in Section 2.4, the threshold is associated with fast convergence. We will show in Section 4.1 that a higher threshold will lead to slower convergence for heterogeneous agents. The detailed discussion on how to choose the threshold is deferred to Remark 4.5.
Theorem 3.1 Proof Sketch   The proof consists of two steps: 1) We reformulate the problem as an MDP where the state space is the Cartesian product of the individual state space, i.e., $\mathcal{S}^{M}$; the action space is the subset of $\{0,1\}^{M}$ that satisfies the budget constraint, and the transition matrix is the product of $P^{m}$ as all agents are independent. 2) We show that the proposed greedy policy $\pi_{t}^{\text{greedy}}$ satisfies the Bellman equation (Eq. (E.21)) for all $t\leq T$ by using backward induction. We prove the induction hypothesis by carefully controlling the difference of the value function (Eq. (E.11)).
4 Learning in BT-CMAB
@@
13:   $\forall a\in\mathcal{A}:\hat{\phi}^{m}(s_{t}^{m},a)=\hat{\phi}^{m}(s_{t}^{m},a)+r^{m}_{t+1}\mathbb{I}\left\{a_{t+1}^{m}=a\right\}$.
14:   $\forall a\in\mathcal{A}:N_{s,m,a}^{(t+1)}=N_{s,m,a}^{(t)}+\mathbb{I}\left\{a_{t+1}^{m}=a\right\}\mathbb{I}\left\{s=s_{t}^{m}\right\}$
15:   $N_{s,m}^{(t+1)}=N_{s,m}^{(t)}+\mathbb{I}\{s=s_{t}^{m}\}$
@@
As we address a thresholding contextual bandit with MDP-governed contexts instead of MAB, Alg. 1 differs from Auer et al. (2002). We tailor $\epsilon_{t}$ to decay with the minimum state occupancy time rather than the current timestep, ensuring sufficient exploration across states.
@@ 4.1 Theoretical Guarantee
We establish that Algorithm 1 achieves a $\mathcal{O}(\sqrt{T})$ regret upper bound under two assumptions. Assump. 4.1 ensures strictly positive transition probabilities (even if arbitrarily small), while Assump 4.2 guarantees the existence of a policy with sufficient good agent-state pairs at each timestep.
@@
Assump. 4.2 serves primarily to simplify our theoretical analysis by ensuring problem instances are well-behaved. This condition guarantees that an oracle policy exists which can spend its full budget of $B$ actions per round without being forced to assign actions with negative expected returns. In practice, this assumption can be satisfied for any problem instance by carefully choosing the threshold parameter $\gamma$; for example, setting $\gamma=$-$\infty$ guarantees the condition is met trivially. Thus, $\gamma$ serves as a tunable parameter to be optimized for overall performance, a point we discuss further in Remark 4.5.
When combined with Assump. 4.1, Assump. 4.2 yields a stronger guarantee (Lemma 4.3) that is inherently assumed by any algorithm required to spend the full budget $B$ at each round. This is common in prior RB algorithms (Wang et al., 2020, 2023b).
Let $\mathcal{G}:=\{(m,s):m\in[M],s\in\mathcal{S},I^{m}(s)\geq\gamma\}$ be the set of good agent-state pairs. Lemma 4.3 ensures that for any state realization $\bm{s}\in\mathcal{S}^{M}$, there are always at least $B$ good agent-state pairs.
@@
Despite its strong guarantee, Lemma 4.3 (proof in Appendix F.1) is derived from two relatively mild assumptions. This result is critical for two reasons. First, it ensures that any algorithm required to spend a full budget of $B$ actions is never forced to assign actions whose incremental reward is below the threshold. Second, it guarantees that while choosing a suboptimal agent may affect future state distributions, this effect does not accumulate: sufficient good agents are always available. Thus, a policy achieving zero per-step regret is attainable, which forms the foundation for our algorithm’s performance.
Let $\Delta_{s}^{m}:=|\gamma-I^{m}(s)|$. Define $(m_{\min},s_{\min})$ as the minimizer of $I^{m}(s)-\gamma$ over $(m,s)\in\mathcal{G}$, and set $\Delta:=\Delta_{s_{\min}}^{m_{\min}}$. Next, we present our main theorem on the regret upper bound, which characterizes the finite-horizon convergence rate of Algorithm 1:
@@
Assume $r_{t}^{m}$ is 1-subgaussian random variable for all $t\in[T]$. Let $D:=\max\left\{\frac{8M}{B},\frac{8M}{M-B},2\right\},p=1,\eta=\Delta$. Under Assump.s 4.2 and 4.1, the regret, $\mathcal{R}(\pi)$, defined by Def. 2.2 with the loss function $l^{m}$ given by Eq. (3), satisfies:
@@
Theorem 4.4 Proof Sketch   Our analysis deviates from standard CMAB analysis because estimating the incremental reward requires pulling suboptimal arms; consequently, we must lower bound the number of times suboptimal arms are pulled, in addition to the usual upper bound. The argument proceeds in three steps. 1) Regret equivalence: we show that regret is equivalent to the expected number of interventions given to bad agents. 2) Bounding the interventions: we prove that this expected number can be bounded either by the overestimation of the incremental reward of bad agents or by the underestimation of that of good agents. We further establish that at each time step, both the number of interventions and non-interventions assigned to agents are lower bounded. 3) Concentration: we apply concentration inequality (Lemma F.2) to obtain the desired result. Importantly, stationarity of the reward is not required here, since the concentration inequality we employ holds without such an assumption.
@@
In our problem, there is an inherent trade-off between satisfying Assump. 4.2 and maximizing cumulative reward. While our reformulation focuses on the threshold identification problem, maximizing cumulative reward remains our ultimate goal. Specifically, a smaller threshold $\gamma$ makes it easier to satisfy Assump. 4.2, as more state-agent pairs qualify as good. However, a low $\gamma$ also makes any arm above the threshold optimal, potentially leading to lower rewards. Empirically, we observe that there exists a universal choice of $\gamma$ which satisfies the Assump. 4.2 that achieves the highest reward as shown in Appendix G.3.
[…]
@@ 5 Experiments
Setup   We set $M=50,\mathcal{S}=\{0,1\}$, and compare the results under three different budget settings: $B=5,10,20$. We use: 1) Synthetic Instances: The expected reward is set to be $R^{m}(s)=s$ and randomly generate 50 instances with different transition matrices, 2) Wireless Communication Dataset (Prieto-Cerdeira et al., 2010): We use the same wireless communication dataset as Wang et al. (2020). Details will be deferred to Section 5.2. We also consider the setting where $\mathcal{S}=[10]$ and randomly generate 50 instances. The results are deferred to Appendix G.2.
In Figure 1, each line is averaged over $50$ instances, each with $13$ repetitions. We consider two reward settings: 1) a noiseless case, and 2) a noisy setting where the rewards at state $s\in\mathcal{S}$ follow a normal distribution with mean $R^{m}(s)$ and variance 1 for the synthetic dataset and 0.1 for the real-world dataset. We note that we did not fine-tune our hyperparameters. Specifically, we universally choose a threshold that satisfies Assump. 4.2. We have also shown that our algorithm is robust across different $\eta$ and we have done ablations on the choice of threshold (Appendix G.3).
5.1 Synthetic Instances
@@
[…]
5.2 Wireless Communication Dataset
This problem (Prieto-Cerdeira et al., 2010) involves digital video broadcasting satellite services to handheld devices via land mobile satellite. In this problem, the environment is modeled as a two-state system (“good” and “bad”), with five different transition models corresponding to different satellite locations. In their original problem, each agent is provided with multiple actions which is modeled as the angle of the satellite. Similar to Wang et al. (2020), we choose two angles as the action: 40 degrees as action 0 and 60 degrees as action 1. The transition probability is chosen according to Table IV in Prieto-Cerdeira et al. (2010). The original reward is dependent on both action and state, but the reward difference between different actions are low (Table III in Prieto-Cerdeira et al. (2010)). Further, the reward difference between the good state and bad state is also small. Therefore, we reconstruct the reward and set the reward for good state as 1 and bad state as 0.9 to mimic the real-world situations where signal is weak. We randomly assign 50 agents with one of the five transition probabilities.
In Figure 2, we observe that in this problem, Our method consistently outperforms all benchmarks within a horizon of 1000 in all budget settings under noiseless and noisy rewards. As shown in Figure 2, Alg. 1 converges to its oracle policy (OracleGreedy) quickly while other online RB algorithms converge to their oracle policies more slowly. This translates to the superior performance of Alg. 1, demonstrating our observation that quick convergence translates to better finite sample performance.
Discussions   We establish greedy policy optimality in finite-horizon RBs and reformulate online RB as a BT-CMAB with a learning algorithm. Our approach leverages the insight that fast convergence to a suboptimal policy can outperform slow convergence to a better one in finite horizons. As the BT-CMAB formulation is novel, our work appropriately focuses on establishing its theoretical and empirical value. Deriving a problem-specific lower bound is a natural and important direction for future work. Further, we plan to analyze the optimality of the oracle policy in multi-state setting.
Checklist
  1. 1.
    For all models and algorithms presented, check if you include:
    1. (a)
      A clear description of the mathematical setting, assumptions, algorithm, and/or model. [Yes]
    2. (b)
      An analysis of the properties and complexity (time, space, sample size) of any algorithm. [Yes]
    3. (c)
      (Optional) Anonymized source code, with specification of all dependencies, including external libraries. [Yes]
  2. 2.
    For any theoretical claim, check if you include:
    1. (a)
      Statements of the full set of assumptions of all theoretical results. [Yes]
    2. (b)
      Complete proofs of all theoretical results. [Yes]
    3. (c)
      Clear explanations of any assumptions. [Yes]
  3. 3.
    For all figures and tables that present empirical results, check if you include:
    1. (a)
      The code, data, and instructions needed to reproduce the main experimental results (either in the supplemental material or as a URL). [Yes]
    2. (b)
      All the training details (e.g., data splits, hyperparameters, how they were chosen). [Yes]
    3. (c)
      A clear definition of the specific measure or statistics and error bars (e.g., with respect to the random seed after running experiments multiple times). [Yes]
    4. (d)
      A description of the computing infrastructure used. (e.g., type of GPUs, internal cluster, or cloud provider). [Yes]
  4. 4.
    If you are using existing assets (e.g., code, data, models) or curating/releasing new assets, check if you include:
    1. (a)
      Citations of the creator If your work uses existing assets. [Not Applicable]
    2. (b)
      The license information of the assets, if applicable. [Not Applicable]
    3. (c)
      New assets either in the supplemental material or as a URL, if applicable. [Not Applicable]
    4. (d)
      Information about consent from data providers/curators. [Not Applicable]
    5. (e)
      Discussion of sensible content if applicable, e.g., personally identifiable information or offensive content. [Not Applicable]
  5. 5.
    If you used crowdsourcing or conducted research with human subjects, check if you include:
    1. (a)
      The full text of instructions given to participants and screenshots. [Not Applicable]
    2. (b)
      Descriptions of potential participant risks, with links to Institutional Review Board (IRB) approvals if applicable. [Not Applicable]
    3. (c)
      The estimated hourly wage paid to participants and the total amount spent on participant compensation. [Not Applicable]
@@ Appendix A Real-World Restless Bandit Examples
We will further use an example to show case how the state, action, reward and budget is defined in practice. SwipeRX is an online platform in Indonesia where pharmacists typically place orders on Mondays. The platform aims to increase engagement and revenue by providing targeted coupons to pharmacists. Since pharmacists operate in different local markets with varying demand patterns and competition from other platforms, we model them as heterogeneous agents, each represented by an MDP.
  • State: The pharmacist’s engagement level with the platform
  • Action: Binary decision of whether to provide a coupon to a given pharmacist
  • Action-dependent State Transitions: Based on whether an action is taken or not, the agent’s engagement evolves accordingly based on environmental factors and platform interactions
  • Reward: Engagement metrics (e.g., time spent on platform, order frequency)
  • Budget: The platform has a budget on the number of coupons available at each day.
@@ Appendix B Whittle’s Index
$Q^{m}_{\lambda}(s,a)+\beta^{*}_{m}$ $=R^{m}(s)-\lambda\mathbb{I}\{a=1\}$
@@
$\beta^{*}_{m}:=\max~\liminf_{T\to\infty}\frac{1}{T}\mathbb{E}\left[\sum_{t=0}^{T}R^{m}\left(s_{t}^{m}\right)-\lambda\mathbb{I}\{a=1\}\mid s_{0}^{m}\right].$
@@
Theorem B.1 (Theorem 2, Weber and Weiss 1990).
@@
$Q^{m}_{\lambda}(s,a)$ $=R^{m}(s)-\lambda\mathbb{I}\{a=1\}$
@@ Appendix E Technical Details In Section 3
$=\frac{I(0)-I(1)}{R(1)-R(0)}\left(F_{\bm{a}_{h+1},h+1}\left((1,0)\right)-F_{\bm{a}_{h+1},h+1}\left((0,0)\right)\right)-\frac{\left(q^{1}_{11}I(0)-q^{1}_{01}I(1)\right)}{R(1)-R(0)}$
$\left(F_{\bm{a}_{h+1},h+1}\left((1,0)\right)-F_{\bm{a}_{h+1},h+1}\left((0,0)\right)\right)$
@@
$=\frac{I(0)-I(1)}{R(1)-R(0)}\left(F_{\bm{a}_{h+1},h+1}\left((1,0)\right)-F_{\bm{a}_{h+1},h+1}\left((0,0)\right)\right)-\frac{\left(q^{1}_{11}I(0)-q^{1}_{01}I(1)\right)}{R(1)-R(0)}$
$\left(F_{\bm{a}_{h+1},h+1}\left((1,0)\right)-F_{\bm{a}_{h+1},h+1}\left((0,0)\right)\right)$